# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 4.3. The $L_{v}^{r}$– estimate for $Q_{0}$
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{q_{0}^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{lq_{0}^{\prime}}}|f_{k}(v^{*})|^{p_{0}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g_{l}(v_{1}^{*})|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{0}^{\prime}}\,d\sigma\,dv_{1}.$
By (2.25) we have $I_{1}(v)\leq{\langle}v{\rangle}^{q_{0}^{\prime}-2}\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}\leq\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}$, since $q_{0}^{\prime}\leq 2$.
For $I_{2}$, we use (2.7) and the substitution $y:=u=v-v_{1}$ to write
$I_{2}(v)$ $=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-R_{\sigma}^{-}(u))|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{0}^{\prime}}\,dv_{1}\,d\sigma$
$=\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-R_{\sigma}^{-}(y))|^{q_{0}}|\widehat{R}_{\sigma}^{-}(y)\cdot\sigma|^{2p_{0}^{\prime}}\,dy\,d\sigma.$
We now use Proposition 2.1 to substitute $\nu=R_{\sigma}^{-}(y)$ and obtain
$I_{2}(v)$ $\lesssim\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|g_{l}(v-\nu)|^{q_{0}}|\widehat{\nu}\cdot\sigma|^{2(p_{0}^{\prime}-1)}\,d\nu\,d\sigma\lesssim\int_{{\mathbb{R}}^{3}}|g_{l}(v-\nu)|^{q_{0}}\,d\nu=\|g_{l}\|_{L_{v}^{q_{0}}}^{q_{0}}.$
Combining (4.29) with the estimates for $I_{1},I_{2}$, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L^{p_{0}}_{v}}^{p_{0}/q_{0}^{\prime}}\|g_{l}\|_{L^{q_{0}}_{v}}^{q_{0}/p_{0}^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\right)^{1/r}.$
$\|$ ${\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}^{r-p_{0}}\|g_{l}\|_{L_{v}^{q_{0}}}^{r-q_{0}}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})|^{p_{0}}|g_{l}(v_{1}^{*})|^{q_{0}}\,d\sigma\,dv_{1}\,dv$
(4.30) $=\|f_{k}\|_{L_{v}^{p_{0}}}^{r-p_{0}}\|g_{l}\|_{L_{v}^{q_{0}}}^{r-q_{0}}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v)|^{p_{0}}|g_{l}(v_{1})|^{q_{0}}\,d\sigma\,dv_{1}\,dv\approx\|f_{k}\|_{L^{p_{0}}_{v}}^{r}\|g_{l}\|_{L^{q_{0}}_{v}}^{r}.$
$\bullet$ Case $p_{0}=1,\,q_{0}=r$ : We use again (4.8) and Hölder’s inequality, to bound instead as follows
@@
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})|^{1/r^{\prime}}\right)\left(|f_{k}(v^{*})|^{1/r}|g_{l}(v_{1}^{*})|\right)$
@@
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
@@
(4.31) $\lesssim\widetilde{I}_{1}^{1/r^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r},$
@@
$\widetilde{I}_{1}(v)=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{r^{\prime}}}{{\langle}v_{1}^{*}{\rangle}^{lr^{\prime}}}|f_{k}(v^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Since $l>2$ and $r\geq 2$, (2.25) implies $\widetilde{I}_{1}(v)\lesssim{\langle}v{\rangle}^{r^{\prime}-2}\|f_{k}\|_{L_{v}^{1}}\leq\|f_{k}\|_{L_{v}^{1}}$. Combining (4.31) with the bound on $\widetilde{I}_{1}$, we obtain
(4.32) ${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L_{v}^{1}}^{1/r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|^{r}\,d\sigma\,dv_{1}\right)^{1/r}.$
$\|$ ${\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}^{r}\lesssim\|f_{k}\|_{L^{1}_{v}}^{r-1}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})||g_{l}(v_{1}^{*})|\,d\sigma\,dv_{1}\,dv$
(4.33) $=\|f_{k}\|_{L^{1}_{v}}^{r-1}\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}|f_{k}(v)||g_{l}(v_{1})|\,d\sigma\,dv_{1}\,dv\approx\|f_{k}\|_{L^{1}_{v}}^{r}\|g_{l}\|_{L^{r}_{v}}^{r}.$
Combining (4.30), (4.33), we deduce
(4.34) $\|{\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L^{r}_{v}}\lesssim\|f_{k}\|_{L^{p_{0}}_{v}}\|g_{l}\|_{L^{q_{0}}_{v}},\quad 1\leq p_{0}\leq 2\leq q_{0}\leq r.$
Conclusion of the proof
Combining (4.5) with (4.27), (4.28), (4.34), estimate (4.26) follows.
4.4. The $L_{v}^{r}$– estimate for $Q_{1}$
Proposition 4.4.
Let $0<\gamma\leq 1$ and $2<r<\infty$. Consider $1<p_{1}<2<q_{1}<r$ and $1\leq p_{2}<2<q_{2}\leq r$ with $\frac{1}{p_{1}}+\frac{1}{q_{1}}=\frac{1}{p_{2}}+\frac{1}{q_{2}}=1+\frac{1}{r}$. Then for any $k\geq l>2$, there holds the estimate
(4.35) $\|{\langle}v{\rangle}^{k}Q_{1}(f,g)\|_{L_{v}^{r}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p_{1}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q_{1}}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q_{2}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p_{2}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$. We use the decomposition (4.14) from Proposition 4.2.
Estimate for $Q_{1}^{sm}$, $Q_{1}^{r}$
The operators $Q_{1}^{sm},Q_{1}^{r}$ can be treated similarly to $Q_{1}^{sm},Q_{1}^{r}$ from Proposition 4.2, we just use Theorem 3.2 instead of Theorem 3.1. Omitting details, we obtain
(4.36) $\|{\langle}v{\rangle}^{k}Q_{1}^{sm}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}},$
(4.37) $\|{\langle}v{\rangle}^{k}Q_{1}^{r}(f,g)\|_{L_{v}^{r}}$ $\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{k}\|_{L_{v}^{q_{1}}}.$
We now estimate the singular operators $Q_{1}^{s_{1}}$ and $Q_{1}^{s_{2}}$.
Estimate for $Q_{1}^{s_{1}}$
We will take different cases for $1<p_{2}<2<q_{2}<r$ and $p_{2}=1,\,q_{2}=r$.
$\bullet$ Case $1<p_{2}<2<q_{2}<r$ : Since $q_{2}>2$, we have $q_{2}^{\prime}<2$, thus
(4.38) $\frac{1}{r}$ $=1-\frac{1}{p_{2}^{\prime}}-\frac{1}{q_{2}^{\prime}}=1-\frac{2}{q_{2}^{\prime}}+\frac{1}{q_{2}^{\prime}}-\frac{1}{p_{2}^{\prime}}<\frac{1}{q_{2}^{\prime}}-\frac{1}{p_{2}^{\prime}},$
Pick $\alpha>1$ with $\frac{\alpha}{r}<\frac{1}{q^{\prime}_{2}}-\frac{1}{p^{\prime}_{1}}$, which is possible to pick due to (4.38). Using (4.19) and Hölder’s inequality, we bound
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|^{q_{2}/p_{2}^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\right)\left(\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q_{2}^{\prime}}|g_{k}(v_{1}^{*})|^{p_{2}/q_{2}^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left(\frac{|f_{l}(v^{*})|^{q_{2}/r}|g_{k}(v_{1}^{*})|^{p_{2}/r}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha/r}}\right)\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.39) $\leq I_{1}^{1/p_{2}^{\prime}}(v)I_{2}^{1/q_{2}^{\prime}}(v)\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}|g_{k}(v_{1}^{*})|^{p_{2}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}\right)^{1/r},$
where
$I_{1}(v)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p_{2}^{\prime}(\frac{1}{q_{2}^{\prime}}-\frac{\alpha}{r})}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$

### v4

@@ 4.2. The $L_{v}^{r}$– estimate
$I_{1}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p^{\prime}(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}|f_{l}(v^{*})|^{q}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv,$
@@
$I_{2}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq^{\prime}}}|g_{k}(v_{1}^{*})|^{p}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv,$
@@
$I_{3}$ $=\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}{\langle}v_{1}^{*}{\rangle}^{1-\alpha}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $q^{\prime}l>2$, we use (2.26) to bound
@@
$I_{2}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\left(\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lq^{\prime}}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\right)\,dv\lesssim\|g_{k}\|_{L^{p}_{v}}^{p}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
For $I_{1}$, since $|v|\lesssim|v_{1}^{*}|$ and $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$, we first bound
$I_{1}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p^{\prime}(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}|f_{l}(v^{*})|^{q}|\phi(v)|^{r^{\prime}}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}\,dv.$
Then, since $p^{\prime}(k-\gamma)>q^{\prime}(k-\gamma)\geq 2$, we argue similarly as in the estimate of $I_{1}$ in the proof of Proposition 4.1, to obtain
$I_{1}$ $\lesssim\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{1-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\,dx\,d\nu\,dv.$
The inner integral can be estimated as follows
$\int_{0}^{1}$ $\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}\,dx\lesssim 1+\int_{1/\sqrt{2}}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}}x\,dx\approx 1+\int_{{\langle}v{\rangle}^{-1}}^{1/2}z^{-p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})}\,dz\lesssim{\langle}v{\rangle}^{p^{\prime}(\frac{1}{q^{\prime}}-\frac{\alpha}{r})-1},$
where we used the substitution $z:=1-x^{2}$, and the fact that $\frac{\alpha}{r}<\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}$ to estimate the integral in $z$.
We obtain
$I_{1}\lesssim\int_{{\mathbb{R}}^{3}}|\phi(v)|^{r^{\prime}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\,d\nu\,dv=\|f_{l}\|_{L^{q}_{v}}^{q}\|\phi\|_{L^{r^{\prime}}_{v}}^{r^{\prime}}.$
Finally, to estimate $I_{3}$, we use the fact that $|v|\leq E^{1/2}$, followed by (2.12) and (2.28) to bound
$I_{3}$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{\frac{1-\alpha}{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2\alpha}}\mathds{1}_{E>16}\mathds{1}_{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}>E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$\approx\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}E^{\frac{1-\alpha}{2}}\frac{|f_{l}(v^{*})|^{q}|g_{k}(v_{1}^{*})|^{p}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\mathds{1}_{E>16}\mathds{1}_{-1\leq\widehat{u}\cdot\sigma<1-2E^{-1/2}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{-1\leq\widehat{u}\cdot\sigma\leq 1-2E^{-1/2}}}{\left(1-\widehat{u}\cdot\sigma\right)^{\alpha}}\,d\sigma\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{-1}^{1-2E^{-1/2}}\frac{1}{\left(1-x\right)^{\alpha}}\,dx\,dv_{1}\,dv$
$=\int_{{\mathbb{R}}^{6}}E^{\frac{1-\alpha}{2}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\mathds{1}_{E>16}\int_{2E^{-1/2}}^{2}z^{-\alpha}\,dz\,dv_{1}\,dv$
$\lesssim\int_{{\mathbb{R}}^{6}}|f_{l}(v)|^{q}|g_{k}(v_{1})|^{p}\,dv_{1}\,dv=\|f_{l}\|_{L_{v}^{q}}^{q}\|g_{k}\|_{L_{v}^{p}}^{p},$
where we used the fact that $\alpha>1$ to estimate the integral in $z$.
Combining (4.16) with the bounds on $I_{1},I_{2},I_{3}$, we deduce
$|I(\phi)|\lesssim\|f_{l}\|_{L^{q}_{v}}\|g_{k}\|_{L^{p}_{v}}\|\phi\|_{L^{r^{\prime}}_{v}}.$
Since $\phi\in L^{r^{\prime}}_{v}$ was chosen arbitrarily, we conclude
(4.17) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(f,g)\|_{L_{v}^{r}}\lesssim\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
Estimate for $Q_{s,2}^{+}$
Again, we argue by duality. Fixing $\phi\in L_{v}^{r^{\prime}}$, we estimate the quantity
$J(\phi):=\int_{{\mathbb{R}}^{3}}{\langle}v{\rangle}^{k+1-\gamma}Q_{s,2}^{+}(v)\phi(v)\,dv.$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
$|J(\phi)|$ $\lesssim\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{6}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{q^{\prime}}-\frac{1}{p^{\prime}}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|^{p/q^{\prime}}|\phi(v)|^{r^{\prime}/q^{\prime}}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p^{\prime}}}\right)$
$\hskip 85.35826pt\times\left(\frac{{\langle}v{\rangle}^{2/p^{\prime}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p^{\prime}}|g_{k}(v_{1}^{*})|^{q/p^{\prime}}|\phi(v)|^{r^{\prime}/p^{\prime}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 85.35826pt\times\left({\langle}v{\rangle}^{1/r}|f_{l}(v^{*})|^{p/r}|g_{k}(v_{1}^{*})|^{q/r}|\right)\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}\,dv$
(4.18) $\leq J_{1}^{1/q^{\prime}}J_{2}^{1/p^{\prime}}J_{3}^{1/r},$